OpenStudy (anonymous):

Solve each of the following problems using a system of equations. Show all of your work and state your solution in a complete sentence. 1.)Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each? 2.)Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770? can someone please help please??

OpenStudy (anonymous):

1. t + b = 56 b = 2t + 8 can you solve this with substitution ?

OpenStudy (anonymous):

just sub in 2t + 8 in for b in the 1st equation to solve for t. Once you have t, you can sub it in to the 2nd equation to find b.

OpenStudy (anonymous):

2. a + s = 150 --> a = 150 - s 45a + 28.5s = 4770 now sub 150 - s in for a into the 2nd equation to find s. Once s is found, sub that back into the first equation to find a.

OpenStudy (anonymous):

I don't know how to do the solving, I am having trouble with that part

OpenStudy (anonymous):

1. t + b = 56 b = 2t + 8 we will sub 2t + 8 in for b in the 1st equation t + (2t + 8) = 56 can you solve this ?

OpenStudy (anonymous):

let me try

OpenStudy (anonymous):

t+10t=56

OpenStudy (anonymous):

no....combine like terms t + 2t + 8 = 56 3t + 8 = 56 you cannot add 2t + 8 because they are not like terms now we will subtract 8 from both sides because we want 2t by itself. 3t + 8 - 8 = 56 - 8 combine like terms 3t = 56 - 8 3t = 48 now we will divide by 3 to get t by itself 3t/3 = 48/3 t = 16 do you understand ?

OpenStudy (anonymous):

Yes. I forgot that

OpenStudy (anonymous):

so now that we know t = 16, we can sub that into the 1st equation: t + b = 56 16 + b = 56 can you solve for b now ?

OpenStudy (anonymous):

just subtract 16 from both sides

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

40

OpenStudy (anonymous):

you got it.....Tony is 16 and Belinda is 40

OpenStudy (anonymous):

you want to work on the next one now ?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

a + s = 150 (adults + students = 150) 45a + 28.5s = 4770 (45 for adult tickets + 28.5 for student tickets = 4770) do you understand how I got the equations ? Because setting up the equation is the hardest part.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now we will isolate a variable in the first equation...it doesn't matter which one, so which one do you want to isolate ? Your choice.....a or s ?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

a + s = 150 --> a = 150 - s understand ?

OpenStudy (anonymous):

I just subtracted s from both sides to get a by itself

OpenStudy (anonymous):

I understand

OpenStudy (anonymous):

now we will sub 150 - s in for a in the 2nd equation... 45a + 28.5s = 4770 45(150 - s) + 28.5s = 4770 do you know how to distribute the 45 through the parenthesis ?

OpenStudy (anonymous):

you * 45 by 150 and s

OpenStudy (anonymous):

yes...but it is a -s not just an s

OpenStudy (anonymous):

so what did you get ?

OpenStudy (anonymous):

i don't understand. i am still working

OpenStudy (anonymous):

you need help ?

OpenStudy (anonymous):

45*150=6750 did i do that right?

OpenStudy (anonymous):

yes....now do 45 * -s

OpenStudy (anonymous):

or you could just leave the - in and do 45 *s

OpenStudy (anonymous):

i need help with this part

OpenStudy (anonymous):

45(150 - s) + 28.5s = 4770 6750 - 45s + 28.5s = 4770 Understand ?

OpenStudy (anonymous):

i don't understand

OpenStudy (anonymous):

which part....the distribution part ?

OpenStudy (anonymous):

the hole thing

OpenStudy (anonymous):

all we are doing is subbing in 150 - s in for a. Do you understand how we got the 150 - s ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

a + s = 150 we want to isolate a variable.....you chose to isolate a, to do this we must subtract s from both sides. a + s - s = 150 - s the s's on the left side cancel out, leaving us with a = 150 - s. Do you understand ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

our 2nd equation is 45a + 28.5s = 4770 we will sub 150 - s in for a 45(150 - s) + 28.5s = 4770 are you with me so far ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i get it now

OpenStudy (anonymous):

now we will distribute... 45(150 - s) + 28.5s = 4770 45 * 150 - 45*s + 28.5s = 4770 6750 - 45s + 28.5s = 4770 now can you combine like terms ?

OpenStudy (anonymous):

150-28.5=121

OpenStudy (anonymous):

no...... 6750 - 16.5s = 4770 what I did was combined my s terms ... -45s + 28.5s = -16.5s Understand ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now we have 6750 - 16.5s = 4770 we want -16.5s on one side and everything else on the other side. So we need to subtract 6750 from both sides. 6750 - 6750 - 16.5s = 4770 - 6750 this cancels out the 6750 on the left side leaving us with : -16.5s = 4770 - 6750 you still with me ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

dont forget, what you do on one side of the equal sign, you have to do on the other side. -16.5s = -1980 now we will divide both sides by -16.5 to get s by itself -16.5s/-16.5 = -1980/-16.5 s = -1980/-16.5 can you find s now ?

OpenStudy (anonymous):

remember a negative divided by a negative is a positive

OpenStudy (anonymous):

120

OpenStudy (anonymous):

so s = 120 now sub 120 in for s in the equation a + s = 150

OpenStudy (anonymous):

30

OpenStudy (anonymous):

yes....a = 30 let me check out answers to see if they are correct...hold on

OpenStudy (anonymous):

okay. thank you for your help!!! :)

OpenStudy (anonymous):

45a + 28.5s = 4770 45(30) + 28.5(120) = 4770 1350 + 3420 = 4770 4770 = 4770 (correct) adult tickets 30 student tickets 120

OpenStudy (anonymous):

glad I could help :)

OpenStudy (anonymous):

:)